OpenStudy (anonymous):

16 283/1000 in simplest form

ganeshie8 (ganeshie8):

\(\large 16 \dfrac{283}{1000}\) like this ?

OpenStudy (anonymous):

yea

ganeshie8 (ganeshie8):

or like this \(\large \dfrac{16283}{1000}\) ?

OpenStudy (anonymous):

the other one

ganeshie8 (ganeshie8):

first one ?

OpenStudy (anonymous):

the first

ganeshie8 (ganeshie8):

okay so its a mixed number

OpenStudy (anonymous):

yea

ganeshie8 (ganeshie8):

use this : \[\large a \dfrac{b}{c} = a + \dfrac{b}{c}\]

ganeshie8 (ganeshie8):

get common denominator and add the fractions

OpenStudy (anonymous):

im only in 2nd grade i have no idea what that means

ganeshie8 (ganeshie8):

how old are you ?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

im 8

ganeshie8 (ganeshie8):

you need to be over 13 to use this site

ganeshie8 (ganeshie8):

seeya after 5 years

OpenStudy (anonymous):

yea i faked age so i could use it

OpenStudy (anonymous):

im lying im in 6th grade im 13

ganeshie8 (ganeshie8):

sure ?

ganeshie8 (ganeshie8):

which lie should i trust more hmm

OpenStudy (anonymous):

i said im in 2nd to make u just plain gimme the answer

OpenStudy (anonymous):

Antidisestablishmentarianism if i where 8 could i speell that

OpenStudy (anonymous):

Antidisestablishmentarianism

OpenStudy (anonymous):

long word

ganeshie8 (ganeshie8):

fair enough :) i can assis u with solving the problem on ur own if u want just want an answer, whats stopping you from using wolfram ?

OpenStudy (anonymous):

You could have copied it.

OpenStudy (anonymous):

Well, you sound ilke you're in second grade and you copy and pasted that.

OpenStudy (anonymous):

*like

OpenStudy (anonymous):

whats copy paste

OpenStudy (anonymous):

nice try...-.-

OpenStudy (anonymous):

if i where 8 would i know what sex is

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

guys im not 8

OpenStudy (anonymous):

Dammit people i'm closing this damn topic!

OpenStudy (anonymous):

????